# MMGRec: Multimodal Generative Recommendation with Transformer Model

### v1

@@ Abstract.
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
With the popularity of multimedia-centric scenarios like social media and micro-video websites, multimodal recommendation systems are gaining widespread attention and adoption. These systems generally follow the embed-and-retrieve paradigm. The paradigm first incorporates multimodal information with collaborative filtering (CF) (Sarwar et al., 2001) to learn representations for users and items in the embedding stage, and then locates the preferred items for users by measuring inner product similarity in the retrieval stage. Relevant research primarily focuses on enhancing representation learning in the embedding stage. For example, ACF (Chen et al., 2017) introduces a hierarchical attention mechanism to select informative modality content, while MMGCN (Wei et al., 2019) incorporates Graph Convolution Network (GCN) to achieve modal information aggregation and propagation.
@@
To address the issues, recent work (Rajput et al., 2023) inherits the Differentiable Search Index algorithm (Tay et al., 2022) and presents a generative paradigm-based recommendation method, which aims to directly generate the list of user-preferred items without similarity calculation. To achieve this goal, two components, including Semantic ID Assignment and Semantic ID Generation, are designed to answer the questions: how to represent an item with a semantic ID and how to predict the semantic IDs of preferred items for a user. Initially, the method assigns each item a unique semantic ID — a sequence of semantic tokens learned from its textual description. Then, a Transformer (Vaswani et al., 2017) model is trained as the recommendation agent to predict the semantic ID of the next item based on the user’s chronological interaction sequence. Each ID token is yielded with autoregressive decoding, and multiple item IDs are generated with beam search (Sutskever et al., 2014). However, for multimodal recommendation, there are still two technical challenges untouched in the two components:
@@
To overcome the challenges, we design a new ID structure — Rec-ID, and present a Multimodal Generative Recommendation (MMGRec) framework for Rec-ID assignment and generation. Specifically, Rec-ID comprises a sequence of semantic tokens followed by a popularity token, leveraging the correlation between item popularity and semantic information. For Rec-ID assignment, we devise a Graph Residual-Quantized Variational AutoEncoder (Graph RQ-VAE) to fuse item multimodal information with CF information and quantize them into a sequence of semantic tokens. Then, we rank the items with identical semantic tokens by popularity and use their ranking index as the last token to avoid the collision problem. After that, for Rec-ID generation, we train a Transformer-based model to directly generate the preferred items’ Rec-IDs based on users’ historical interactions. To capture item position information in historical interaction sequences, we design a relation-aware self-attention mechanism for Transformer. By extending the original self-attention, this mechanism constructs user-specific relations among items under the supervision of user parameters. We conduct extensive experiments on three public datasets to demonstrate the rationality and effectiveness of the proposed MMGRec.
@@
  • Technically, we design a multimodal information quantization algorithm Graph RQ-VAE for Rec-ID assignment, and a relation-aware self-attention mechanism within Transformer for Rec-ID generation.
  • We conduct empirical studies on three real-world datasets. Extensive results demonstrate that MMGRec achieves state-of-the-art performance with promising inference efficiency.
@@ 2. Preliminary
where $\mathbf{f}^{t}$ is the item’s textual feature converted from descriptions using Sentence-T5 (Ni et al., 2022). $\phi(\cdot)$ indicates the vector quantization algorithm (Zeghidour et al., 2021) converting $\mathbf{f}^{t}$ into a tuple of $M-1$ semantic tokens.
@@
With the obtained semantic ID, a Transformer-based sequence-to-sequence model is trained on the sequences of semantic IDs associated with items from a user’s chronological interactions. The objective is to predict the semantic ID of the next item after the sequence. This model autoregressively decodes the tokens of semantic ID identifying the next item, thereby qualifying as a generative recommendation.
@@ 3. Method
In this section, we first elaborate on our designed MMGRec model according to its two components — Rec-ID assignment and Rec-ID generation. After that, we detail the model training and inference, followed by the discussion regarding the time complexity.
@@ 3.2.1. Graph RQ-VAE
(4) $\mathbf{h}_{i}^{(0)}=\begin{bmatrix}\mathbf{h}_{i}^{m}\,\mathbf{h}_{i}^{c}\end{bmatrix},\,\mathbf{h}_{i}^{m}=\begin{bmatrix}\mathbf{f}_{i}^{v}\,\mathbf{f}_{i}^{a}\,\mathbf{f}_{i}^{t}\end{bmatrix}\mathbf{W},$
where $\mathbf{h}_{i}^{m}\in\mathbb{R}^{D}$ represents item $i$’s multimodal embedding learned by concatenating its visual, acoustic, and textual features, i.e., $\mathbf{f}_{i}^{v}\in\mathbb{R}^{D^{v}}$, $\mathbf{f}_{i}^{a}\in\mathbb{R}^{D^{a}}$, and $\mathbf{f}_{i}^{t}\in\mathbb{R}^{D^{t}}$. $D^{v}$, $D^{a}$, and $D^{t}$ denote the dimensions of these modality features. $\mathbf{W}\in\mathbb{R}^{(D^{v}+D^{a}+D^{t})\times D}$ is a parameter matrix that maps multimodal features to the $D$-dimensional embedding space. Combining this multimodal embedding with a randomly initialized CF embedding $\mathbf{h}_{i}^{c}\in\mathbb{R}^{D}$ yields the item representation $\mathbf{h}_{i}^{(0)}\in\mathbb{R}^{2D}$. For performing graph convolution operations on the bipartite graph, user representations are randomly initialized, such as $\mathbf{h}_{u}^{(0)}\in\mathbb{R}^{2D}$ for user $u$.
Treating them as node embeddings at the 00-th layer, item $i$’s representation at the $n$-th GCN layer is obtained as:
(5) $\mathbf{h}_{i}^{(n)}=\text{LeakyReLU}(\mathbf{h}_{i}^{(n-1)}\mathbf{W}_{1}+\frac{1}{|\mathcal{N}_{i}|}\sum_{u\in\mathcal{N}_{i}}\mathbf{h}_{u}^{(n-1)}\mathbf{W}_{2}),$
where $\mathcal{N}_{i}$ is the neighborhood set of item $i$ (i.e., users whom item $i$ directly interacted with), and LeakyReLU is the activation function. Analogously, the representation $\mathbf{h}_{u}^{(n)}$ for user $u$ is obtained by information propagation from its neighbor items. After stacked GCN layers, final representations are denoted as $\mathbf{h}_{u}\in\mathbb{R}^{D}$ and $\mathbf{h}_{i}\in\mathbb{R}^{D}$. Note that we utilize the concise graph convolution operation (Hamilton et al., 2017), leaving complicated choices (Liu et al., 2023) for further exploration.
@@
(7) $c_{i,1}=\text{argmin}_{l}||\mathbf{r}_{i,1}-\mathbf{b}_{l}||_{2}^{2},$
where $c_{i,1}$ is the index of the closest embedding, i.e., the first codeword. Recursively, for subsequent level $m$, the residual is defined as $\mathbf{r}_{i,m}=\mathbf{r}_{i,m-1}-\mathbf{b}_{c_{i,m-1}}$, and the codeword $c_{i,m}$ is similarly computed using the level-specific codebook $\mathbf{B}_{m}$. By iteratively performing $M-1$ operations, a sequence of codewords is generated as follows:
@@
(10) $\begin{split}\mathcal{L}=\mathcal{L}_{\text{bpr}}&+\mathcal{L}_{\text{rqvae}},\quad\mathcal{L}_{\text{bpr}}=\sum_{(u,i,j)\in\mathcal{R}}-\text{ln}~{}\sigma(\mathbf{h}_{u}\mathbf{h}_{i}^{\top}-\mathbf{h}_{u}\mathbf{h}_{j}^{\top}),\\ \mathcal{L}_{\text{rqvae}}=&\sum_{i=1}^{I}||\mathbf{h}_{i}-\widehat{\mathbf{h}}_{i}||^{2}+\sum_{m=1}^{M-1}||\text{sg}(\mathbf{r}_{i,m})-\mathbf{b}_{c_{i,m}}||^{2}\\ &+\beta||\mathbf{r}_{i,m}-\text{sg}(\mathbf{b}_{c_{i,m}})||^{2}.\end{split}$
Here $\mathcal{L}_{\text{bpr}}$ is the pairwise BPR (Rendle et al., 2009) loss to learn user and item representations. =$\mathcal{R}=\{(u,i,j)|(u,i)\in\mathcal{R}^{+},(u,j)\in\mathcal{R}^{-}\}$, where $\mathcal{R}^{+}$ indicates the observed interactions, and $\mathcal{R}^{-}$ indicates the unobserved interactions. $\sigma(\cdot)$ is the Sigmoid function. The loss function $\mathcal{L}_{\text{rqvae}}$ facilitates the simultaneous training of the encoder-decoder and codebooks in RQ-VAE. $I$ denotes the number of items. We alternatively optimize one of two sub-losses to learn model parameters.
@@ 3.3.1. Transformer Input Embedding
Taking user $u$ as an example, we gather her/his interacted items’ atomic IDs together as the Transformer’s input, and utilize the learned item representations for input embedding, formally,
(13) $\begin{bmatrix}\mathbf{e}_{1},\cdots,\mathbf{e}_{N}\end{bmatrix}=\begin{bmatrix}\mathbf{h}_{i_{1}},\cdots,\mathbf{h}_{i_{N}}\end{bmatrix},$
where $\mathbf{e}_{i}\in\mathbb{R}^{D}$ denotes the embedding of the corresponding item in the sequence. It is worth noting that the input sequence length is uniformly set to $N$, filling the missing positions with padding token 00.
@@ 3.3.2. Relation-Aware Self-Attention
After obtaining the encoded vectors, we conduct the multi-head attention operations on the input sequence. In each head, we map the input element embeddings $\begin{bmatrix}\mathbf{e}_{1},\cdots,\mathbf{e}_{N}\end{bmatrix}$ into query, key, and value spaces, and compute their new embeddings $\begin{bmatrix}\mathbf{x}_{1},\cdots,\mathbf{x}_{N}\end{bmatrix}$ where $\mathbf{x}_{i}\in\mathbb{R}^{D}$:
@@
Overall, we use the relation-aware self-attention sublayer followed by a feed-forward sublayer as the Transformer encoder layer. Residual connections surround each sublayer, followed by layer normalization. As for the Transformer decoder, we maintain the default decoder layers and positional encoding, as the Transformer’s output sequence is inherently ordered.
@@ 3.4. Training & Inference
During model training, each observed interaction ($u,i$) forms a training sample. We input the atomic ID-based interaction sequence of user $u$ into the Transformer encoder. To ensure generative recommendation, we deliberately exclude the ground-truth item $i$ from the input sequence. Model parameters are learned by optimizing cross-entropy loss between Transformer decoder output and item $i$’s Rec-ID. This loss function is commonly used for Transformer, and the indices of Graph RQ-VAE codebooks serve as vocabulary.
In inference with a well-trained Transformer, results are output step by step, where each step’s conditional probability distribution depends on the already generated results. To maximize the product of conditional probabilities for the entire output, we incorporate beam search (Sutskever et al., 2014) to preserve $K$ sequences with the currently highest probabilities at each time step. The beam number is configured to $K$ for top-$K$ recommendation. Through $M$ steps of autoregressive decoding, beam search can ultimately generate $K$ complete Rec-IDs. To prevent generating invalid Rec-IDs, we use constrained beam search (De Cao et al., 2020) to limit the current token range based on prefixes.

### v2

@@ Abstract.
doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Information systems Recommender systemsccs: Information systems Multimedia and multimodal retrievalccs: Computing methodologies Neural networks
@@ 1. Introduction
With the popularity of multimedia-centric scenarios like social media and micro-video websites, multimodal recommendation systems are gaining widespread attention and adoption. These systems generally follow the embed-and-retrieve paradigm. The paradigm first incorporates multimodal information with collaborative filtering (CF) (Sarwar et al., 2001) to learn representations for users and items in the embedding stage, and then locates the preferred items for users by measuring inner product similarity in the retrieval stage. Relevant research primarily focuses on enhancing representation learning in the embedding stage (Liu et al., 2022; De Divitiis et al., 2023). For example, ACF (Chen et al., 2017) introduces a hierarchical attention mechanism to select informative modality content, while MMGCN (Wei et al., 2019) incorporates Graph Convolution Network (GCN) to achieve modal information aggregation and propagation.
@@
To address these issues, recent work TIGER (Rajput et al., 2023) introduces semantic IDs by inheriting the Differentiable Search Index algorithm (Tay et al., 2022) and proposes a generative paradigm-based recommendation framework. This framework aims to directly generate a list of user-preferred items without relying on similarity computation. To achieve this goal, two components, including Semantic ID Assignment and Semantic ID Generation, are designed to answer the questions: how to represent an item with a semantic ID and how to predict the semantic IDs of preferred items for a user. Initially, the method assigns each item a unique semantic ID — a sequence of semantic tokens learned from its textual description. Then, a Transformer (Vaswani et al., 2017) model is trained as the recommendation agent to predict the semantic ID of the next item based on the user’s chronological interaction sequence. Each ID token is yielded with autoregressive decoding, and multiple item IDs are generated with beam search (Sutskever et al., 2014). However, for multimodal recommendation, there are still two technical challenges untouched in the two components:
@@
To overcome the challenges, we design a new ID structure — Rec-ID, and present a Multimodal Generative Recommendation (MMGRec) framework for Rec-ID assignment and generation. Specifically, Rec-ID comprises a sequence of semantic tokens followed by a popularity token, leveraging the correlation between item popularity and semantic information. For Rec-ID assignment, building upon the RQ-VAE adopted in TIGER, we propose a Graph Residual-Quantized Variational AutoEncoder (Graph RQ-VAE) that incorporates a graph network to fuse item multimodal information with collaborative filtering signals, and quantizes the resulting representations into a sequence of semantic tokens. Then, we rank the items with identical semantic tokens by popularity and use their ranking index as the last token to avoid the collision problem. After that, for Rec-ID generation, we train a Transformer-based model to directly generate the preferred items’ Rec-IDs based on users’ historical interactions. To capture item position information in historical interaction sequences, we design a relation-aware self-attention mechanism for Transformer. By extending the original self-attention mechanism, this approach models user-specific relationships among items through personalized query, key, and value projection matrix parameters. We conduct extensive experiments on three public datasets to demonstrate the rationality and effectiveness of the proposed MMGRec.
@@
  • Technically, we propose a multimodal information quantization framework, namely Graph RQ-VAE, which incorporates a popularity token for Rec-ID assignment. We further design a relation-aware self-attention mechanism within the Transformer architecture for Rec-ID generation.
  • We conduct empirical studies on three real-world datasets. Extensive results demonstrate that MMGRec achieves state-of-the-art performance with promising inference efficiency. The codes are released as a side contribution [^1] https://github.com/hanliu95/MMGRec..
2. Related Work
2.1. Multimodal Recommendation
Research on multimodal recommendation aims to enhance recommender systems by leveraging multimodal information (Liu et al., 2023a). Existing methodologies typically employ a combination of CF and content-based recommendation techniques (Liu et al., 2023b). Early studies aim to integrate multimodal content for more rational representation learning. For example, the pioneering work VBPR (He and McAuley, 2016) maps visual information to the representation space and concatenates it with CF representations. Subsequent work like ACF (Chen et al., 2017) and UVCAN (Liu et al., 2019) adaptively select multimodal features using different levels of attention mechanisms to learn representations of users and items. In recent years, the research trend turns to exploring the relationship between modal information and users/items. For instance, the representative work MMGCN (Wei et al., 2019) introduces a modality-aware GCN into multimodal recommendation tasks to aggregate and propagate multimodal information on the user-item graph. HHFAN (Cai et al., 2021) utilizes a modality-aware heterogeneous information graph to explore more complex relationships among users, micro-videos, and multimodal information, thereby improving recommendation performance.
2.2. Transformer-Based Recommendation Model
2.3. LLM-Based Recommendation Systems
In recent years, recommendation systems have increasingly incorporated Large Language Models (LLMs) to exploit their rich knowledge and strong reasoning abilities, mainly in two paradigms: LLM-augmented and LLM-based recommendation systems (Wu et al., 2024).
LLM-augmented recommendation systems typically use LLM-generated embeddings as semantic representations (Hou et al., 2023) or high-level feature extractors (Xi et al., 2024), while still relying on traditional models for final decision making, which limits the use of LLMs’ reasoning ability (Hou et al., 2022). In contrast, LLM-based recommenders directly perform recommendation with LLMs. Early work exploited in-context learning for zero- or few-shot recommendation (Dai et al., 2023), but untuned models suffer from weak instruction following and limited domain knowledge. Recent studies, therefore, adopt supervised fine-tuning with historical interaction data to improve recommendation performance (Bao et al., 2023; Lin et al., 2024).
Our approach is conceptually closer to the second line of work, but differs in that we adopt a from-scratch training paradigm. Compared with fine-tuning large pre-trained models, fully training a small generative recommender offers better efficiency, stronger domain adaptation, and clearer controllability, making it more suitable for scalable and resource-constrained recommendation scenarios.
@@ 3. Preliminary
where $\mathbf{f}^{t}$ is the item’s textual feature converted from descriptions using Sentence-T5 (Ni et al., 2022). $\phi(\cdot)$ indicates the vector quantization algorithm (Zeghidour et al., 2021) converting $\mathbf{f}^{t}$ into a tuple of M1M-1 semantic tokens.
@@
With the obtained semantic ID, a Transformer-based sequence-to-sequence model is trained on the sequences of semantic IDs corresponding to the items from a user’s chronological interactions. The objective is to predict the semantic ID of the next item after the sequence. This model autoregressively decodes the tokens of semantic ID identifying the next item, thereby qualifying as a generative recommendation.
@@ 4. Method
In this section, we first elaborate on our designed MMGRec model, which consists of two stages: Rec-ID assignment and Rec-ID generation. After that, we detail the model training and inference, followed by a discussion regarding the time complexity.
@@ 4.2.1. Graph RQ-VAE
(4) $\mathbf{h}_{i}^{(0)}=\mathbf{h}_{i}^{m}\oplus\mathbf{h}_{i}^{c},\quad\mathbf{h}_{i}^{m}=(\mathbf{f}_{i}^{v}\oplus\mathbf{f}_{i}^{a}\oplus\mathbf{f}_{i}^{t})\mathbf{W},$
where $\oplus$ denotes the vector concatenation operation. $\mathbf{h}_{i}^{m}\in\mathbb{R}^{D}$ represents item $i$’s multimodal embedding learned by concatenating its visual, acoustic, and textual features, i.e., $\mathbf{f}_{i}^{v}\in\mathbb{R}^{D^{v}}$, $\mathbf{f}_{i}^{a}\in\mathbb{R}^{D^{a}}$, and $\mathbf{f}_{i}^{t}\in\mathbb{R}^{D^{t}}$. $D^{v}$, $D^{a}$, and $D^{t}$ denote the dimensions of these modality features. $\mathbf{W}\in\mathbb{R}^{(D^{v}+D^{a}+D^{t})\times D}$ is a parameter matrix that maps multimodal features to the $D$-dimensional embedding space. Combining this multimodal embedding with a randomly initialized CF embedding $\mathbf{h}_{i}^{c}\in\mathbb{R}^{D}$ yields the item representation $\mathbf{h}_{i}^{(0)}\in\mathbb{R}^{2D}$. For performing graph convolution operations on the bipartite graph, user representations are randomly initialized (Lei et al., 2023), such as $\mathbf{h}_{u}^{(0)}\in\mathbb{R}^{2D}$ for user $u$.
Treating them as node embeddings at the 0-th layer, item $i$’s representation at the $y$-th GCN layer is obtained as:
(5) $\mathbf{h}_{i}^{(y)}=\text{LeakyReLU}(\mathbf{h}_{i}^{(y-1)}\mathbf{W}_{1}+\frac{1}{|\mathcal{N}_{i}|}\sum_{u\in\mathcal{N}_{i}}\mathbf{h}_{u}^{(y-1)}\mathbf{W}_{2}),$
where $\mathcal{N}_{i}$ is the neighborhood set of item $i$ (i.e., users whom item $i$ directly interacted with), and LeakyReLU is the activation function. Analogously, the representation $\mathbf{h}_{u}^{(y)}$ for user $u$ is obtained by information propagation from its neighbor items. After stacked GCN layers, final representations are denoted as $\mathbf{h}_{u}\in\mathbb{R}^{D}$ and $\mathbf{h}_{i}\in\mathbb{R}^{D}$. Note that we utilize the concise graph convolution operation (Hamilton et al., 2017), leaving complicated choices (Liu et al., 2023b) for further exploration.
@@
[…]
where $c_{i,1}$ is the index of the closest embedding, i.e., the first codeword. Recursively, for subsequent level $m$, the residual is defined as $\mathbf{r}_{i,m}=\mathbf{r}_{i,m-1}-\mathbf{b}_{c_{i,m-1}}$, and the codeword $c_{i,m}$ is similarly computed using the level-specific codebook $\mathbf{B}_{m}$. By iteratively performing M1M-1 operations, a sequence of codewords is generated as follows:
@@
(10) $\begin{split}&\mathcal{L}=\mathcal{L}_{\text{bpr}}+\mathcal{L}_{\text{rqvae}},\\ \mathcal{L}_{\text{bpr}}=&\sum_{(u,i,j)\in\mathcal{R}}-\text{ln}~\sigma(\mathbf{h}_{u}\mathbf{h}_{i}^{\top}-\mathbf{h}_{u}\mathbf{h}_{j}^{\top}),\\ \mathcal{L}_{\text{rqvae}}=\sum_{i=1}^{I}\big(||\mathbf{h}_{i}-\widehat{\mathbf{h}}_{i}||^{2}&+\sum_{m=1}^{M-1}(||\text{sg}(\mathbf{r}_{i,m})-\mathbf{b}_{c_{i,m}}||^{2}+\beta||\mathbf{r}_{i,m}-\text{sg}(\mathbf{b}_{c_{i,m}})||^{2})\big).\end{split}$
Here $\mathcal{L}_{\text{bpr}}$ is the pairwise BPR (Rendle et al., 2009) loss to learn user and item representations. $\mathcal{R}=\{(u,i,j)|(u,i)\in\mathcal{R}^{+},(u,j)\in\mathcal{R}^{-}\}$, where $\mathcal{R}^{+}$ indicates the observed interactions, and $\mathcal{R}^{-}$ indicates the unobserved interactions. $\sigma(\cdot)$ is the Sigmoid function. The loss function $\mathcal{L}_{\text{rqvae}}$ facilitates the simultaneous training of the encoder-decoder and codebooks in RQ-VAE. $I$ denotes the number of items. We alternatively optimize one of two sub-losses to learn model parameters.
@@ 4.3.1. Transformer Input Embedding
Taking user $u$ as an example, we gather her/his interacted items’ atomic IDs together as the Transformer’s input sequence, e.g., $(i_{1},\cdots,i_{N})$. Then, we utilize the learned item representations for input embedding to convert the input tokens to vectors, formally,
(13) $\begin{bmatrix}\mathbf{e}_{1},\cdots,\mathbf{e}_{N}\end{bmatrix}=\begin{bmatrix}\mathbf{h}_{i_{1}},\cdots,\mathbf{h}_{i_{N}}\end{bmatrix}=\text{Emb}(i_{1},\cdots,i_{N}),$
where $\mathbf{e}_{i}\in\mathbb{R}^{D}$ denotes the embedding of the corresponding item in the sequence. It is worth noting that the input sequence length is uniformly set to $N$, filling the missing positions with padding token 0.
@@ 4.3.2. Relation-Aware Self-Attention
After obtaining the encoded vectors, we conduct the multi-head attention operations on the input sequence. In each head, we map the input element embeddings $\mathbf{E}=\begin{bmatrix}\mathbf{e}_{1},\cdots,\mathbf{e}_{N}\end{bmatrix}$ into query, key, and value spaces, and compute their new embeddings $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1},\cdots,\mathbf{x}_{N}\end{bmatrix}$ where $\mathbf{x}_{i}\in\mathbb{R}^{D}$:
@@
In our overall architecture, we employ a Transformer encoder with relation-aware self-attention. The decoder follows the standard Transformer design for autoregressive Rec-ID generation.
@@ 4.4. Training & Inference
During training, each interaction ($u,i$) is treated as a training instance. The encoder takes the atomic ID-based interaction sequence of user $u$ as input, excluding the target item $i$. The decoder is trained to autoregressively generate the Rec-ID of item $i$ using a cross-entropy loss, where the Graph RQ-VAE codebook indices serve as the output vocabulary. During inference, the trained Transformer generates Rec-IDs in an autoregressive manner. We apply beam search (Sutskever et al., 2014) to maintain the top-$K$ candidate sequences at each decoding step, yielding $K$ complete Rec-IDs after $M$ steps for top-$K$ recommendation. To avoid invalid Rec-ID generation, we further adopt constrained beam search (De Cao et al., 2020), where the candidate token set at each step is restricted based on valid prefixes.